OpenStudy (anonymous):

Tom buys a new phone service. It costs $20 to install, and then costs $5 a month. Write an equation to show the cost of the phone (y) for any number of months (x).

OpenStudy (anonymous):

okay: let y equal the cost. and x equal months. so: $20 is a one time fee. $5 is price for each month therefore: y=5x+20

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can you help me with another one

OpenStudy (anonymous):

yes but you need to open a new question. for this question you can look at my comment and mark "best response" then close the question and type a new one. in the question or in a comment, put @montelbaun. and it will send me a link to your question

OpenStudy (anonymous):

For the straight line y= \[\frac{ 1 }{ 2 }x\]-4 , what are the slope and y-intercept?

OpenStudy (anonymous):

okay ill answer it on here but next time open a new question :)

OpenStudy (anonymous):

are you familiar with the general slope-intecept formula: y=mx+b

OpenStudy (anonymous):

yeah kinda not really.

OpenStudy (anonymous):

okay well that is what is used to solve this easily. in y=mx+b: m=slope b=y-intercept. therefore in y=\[\frac{ 1 }{ 2 }x-4\] m=\[\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

and the y-intercept is -4 because: